OpenStudy (anonymous):

Write this expression in exponential form: Sqrt x^3

OpenStudy (anonymous):

\[x^{\frac{3}{2}}\]

OpenStudy (anonymous):

satellite is correct, the answer is obtained as the sqrt is the same as 1/2 so x^3/1 times 1/2 is x^3/2

OpenStudy (anonymous):

Thank you so much! Could you help me with this next problem?

OpenStudy (anonymous):

3 Sqrt of 2y^2

OpenStudy (anonymous):

√(x^3 ) √ in power/exponential form is 1/2 x^(3/1 times 1/2) = x^(3/2)

OpenStudy (anonymous):

that's to put it in a more simple way for the first problem.

OpenStudy (anonymous):